OpenStudy (anonymous):

Simplify x^-6 . x^4 write your answer with a positive exponent only

OpenStudy (luigi0210):

Rules to know: \[\Huge k^{-a}=\frac{1}{k^a}\] \[\Huge\frac{k^a}{k^b}=k^{a-b}\]

OpenStudy (anonymous):

can you do it?

OpenStudy (anonymous):

referring to that formula, pretend a=-6 and b=4

OpenStudy (anonymous):

i got it thanks